3. If the matrices A and B are given by A= and B =
a. Calculate 3A -2B+7I
b. 3A + 4X = B is an equation for an unknown matrix X. Find the matrix X
c. Calculate the matrix product AB
4. Solve the following set of equations by Gaussian elimination:
3x + 2y + z =2
4x + 2y + 2z =8
x – y + z = 8
1
Expert's answer
2020-04-29T18:32:35-0400

let"A=\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix}" "B=\begin{bmatrix}
 i & j\\
 k & l
\end{bmatrix}"

then

a. Calculate 3A -2B+7I

"3A=\begin{bmatrix}
 3a & 3b \\
 3c & 3d
\end{bmatrix}\\2B=\begin{bmatrix}
 2i & 2j\\
 2k & 2l
\end{bmatrix}\\7I=\begin{bmatrix}
 7& 0 \\
 0& 7
\end{bmatrix}\\3A-2B+7I=\begin{bmatrix}
 3a & 3b \\
 3c & 3d
\end{bmatrix}-\begin{bmatrix}
 2i & 2j\\
 2k & 2l
\end{bmatrix}+\begin{bmatrix}
 7 &0\\
 0& 7
\end{bmatrix}"


"3A-2B+7I=\begin{bmatrix}
 3a-2i+7 & 3b-2j \\
 3c-2k & 3d-2l+7
\end{bmatrix}"


b.) 3A + 4X = B is an equation for an unknown matrix X. Find the matrix X

let "X=\begin{bmatrix}
 u&v\\
 w& z
\end{bmatrix}"

then


"\begin{bmatrix}
 3a & 3b \\
 3c & 3d
\end{bmatrix}+\begin{bmatrix}
 4u& 4v \\
 4w & 4z
\end{bmatrix}=\begin{bmatrix}
 i & j \\
 k & l
\end{bmatrix}"


forming system of linear equations

"3a+4u=i\\3b+4v=j\\3c+4w=k\\3d+4z=l"

solving for u, v, w and z

"u=\frac{i-3a}{4}\\v=\frac{j-3b}{4}\\w=\frac{k-3c}{4}\\z=\frac{l-3d}{4}"

hence

"X=\begin{bmatrix}
 \frac{i-3a}{4}& \frac{j-3b}{4} \\
 \frac{k-3c}{4}& \frac{l-3d}{4}
\end{bmatrix}"


c.) Calculate the matrix product AB

"A=\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix};B=\begin{bmatrix}
 i& j\\
 k & l
\end{bmatrix}"

then

"\\\\\\AB=\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix}\begin{bmatrix}
 i& j\\
 k & l
\end{bmatrix}\\\\
AB=\begin{bmatrix}
 ai+bk& aj+bl\\
 ci+dk &cj+dl
\end{bmatrix}"


4.) Solve the following set of equations by Gaussian elimination:

3x + 2y + z =2

4x + 2y + 2z =8

x – y + z = 8

extract augmented matrix

"\begin{bmatrix}
 3 & 2&1&2 \\
 4& 2&2&8\\
1&-1&1&8
\end{bmatrix}\\R_{1} to R _{3}\begin{bmatrix}
 3 & 2&1&2 \\
 4& 2&2&8\\
1&-1&1&8
\end{bmatrix}
\\\begin{bmatrix}
 1 & -1&1&8 \\
 4& 2&2&8\\
3&2&1&2
\end{bmatrix}"

divide R2 by 2

"\begin{bmatrix}
 1 & -1&1&8 \\
 2& 1&1&4\\
3&2&1&2
\end{bmatrix}"

using row1 to reduce row2 and row3

"\begin{bmatrix}
 1 & -1&1&8 \\
 0& 3&-1&-12\\
0&5&-2&-22
\end{bmatrix}"

using row2 to reduce row3 and keep row1

"\begin{bmatrix}
 1 & 0&2&12 \\
 0& 3&-1&-12\\
0&0&-1&-6
\end{bmatrix}"

using row3 to reduce row1and keep row2

"\begin{bmatrix}
 1 & 0&0&0 \\
 0& 3&0&-6\\
0&0&-1&-6
\end{bmatrix}"


dividing R2 by 3 and R3 by -1

"\begin{bmatrix}
 1 &0&0&0 \\
 0& 1&0&-2\\
0&0&1&6
\end{bmatrix}"

hence

"x=0\\y=-2\\z=6"